AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 11 of 33  ELOG logo
ID Date Author Subject
  462   Wed Jan 10 10:22:41 2024 CB, JMSome FEEs mounted - issues with Det#2 (bottom left)

Started mounting some FEEs. We ran into some mechanical issues due to the new ERNI connectors on the new flexi-rigid adaptor PCBs. Davide skimmed 6.5 mm off the edge with the groove (the groove is gone). This is suffiicent to support the flexi-rigid and in spite of our earlier concerns, the flexi-rigid can be mated/un-mated with the FEEs without major alignment issues.

We still have issues for the bottom-right detector, where the FEE holder will just not fit - will probably figure something out with Davide.

 

We observe a peculiar issue with the bottom-left detector. At 150V, the leakage current for this detector as tested with just the adaptor cards on the feedthroughs that carry the bias (no FEEs) is of the order of 3 uA.

We mounted all four flexi-rigid PCBs - still 3 uA.

We mounted the FEE that slots into the bottom-most feedthrough (which corresponds to the n-n bias). Still 3 uA.

We mounted the FEE that slots in to the second bottom feedthrough (not corresponding to bias). 15 uA! Upon lowerint the bias, the leakage current goes to zero around 30V as opposed to 0V like all the other detectors, and like this detector when this particular FEE is not mounted.

Tests were performed exchanging all PCBs, and all FEEs. The same behaviour is observed. It is not an issue with the FEE or the PCB. We suspect some cross-contact between detector bias and some strip that feeds into the second FEE? In any case, it is most likely a vacuum side issue - nothing we can do now.

 

The detector can be biased to 80V or so without major issue, and it appears this behaviour improves over time, i.e. leaving the bias on reduces the leakage current dramatically so it may even be possible to reach 150V.

See plots attached. Note in 2nd plot the plotting script has some issues - detector names are correct until the bias in Ch#2 is brough to 150V, then the colours are incorrect. Only Ch#2 was being changed. Note the drop in leakage current by a factor 2 just increasing the bias to 150V for a few seconds and returning down.

Attachment 1: Screenshot_2024-01-10_11-59-39.png
Screenshot_2024-01-10_11-59-39.png
Attachment 2: Screenshot_2024-01-10_12-56-36.png
Screenshot_2024-01-10_12-56-36.png
  461   Mon Jan 8 19:29:57 2024 CB, JMManifold repaired + fee assembly in bench

Davide extracted the stuck blind plugs and re-threaded all missing ports. Completed manifold assembly, except for main inlet/outlet - left for plumber.

Cut 8x2 2m hoses and assembled with Swagelok quikfit and teflon ferrules.

We're missing back ferrules in brass for the repaired fees. It should be possible to buy them from GSI.

 

Attempted assembly of fee, Flexi rigid and black delrin support on bench. Davide shimmed the black delrin supports by 6.5 mm from ridge. They now fit length wise on new adaptor cards. Unfortunately, the decision to flip the connector means the Flexi rigid is no longer at the correct height for the erni connector in the fee. See attached.

Inserting the Flexi rigid without the black support is too dangerous, and defeats the point of having rails. Will probably mount Flexi rigids tomorrow on black support, not screw into the fee holder, and align by eye. Tried once and it works - see last two attachments - but it will have unclear impact on the ease to unmount the fee without damaging the Flexi rigid and will make mounting fee holder on flanges harder. No alternative solutions in current timeframe.

Attachment 1: 20240108_164756.jpg
20240108_164756.jpg
Attachment 2: 20240108_163218.jpg
20240108_163218.jpg
Attachment 3: 20240108_163916.jpg
20240108_163916.jpg
Attachment 4: 20240108_164010.jpg
20240108_164010.jpg
  460   Mon Jan 8 19:27:36 2024 CB, JMLeak current tests overnight

Left bias in overnight. Leakage current decreased and stabilised. See attached. Bias off as we prepare to mount the fees.

Attachment 1: Screenshot_2024-01-08_09-56-20.png
Screenshot_2024-01-08_09-56-20.png
  459   Sun Jan 7 15:38:57 2024 CB, JMManifolds mostly assembled + bias overnight check

Assembled most of the connectors on the new (green) manifolds. Two blanking pulgs stuck when mounting - cannot remove them or tighten them. Swarf? Will ask Davide's help tomorrow. Cannot use the manifold until this is fixed.

Checked FEE assembly in cassette. Need to shorten black Delrin supports by ~6 mm. Holes should end up in the correct position thanks to rails being effectively longer (same length, different hole position on FEE cassette). Will ask Davide to skim one if possible. If not, will have to mount without supports.

Finally managed to get the CAEN program working to control N1419 remotely from Baking Pi. Plotting bias and leakage current overnight to test stability. See attached. IE514 sensor off for these tests.

Attachment 1: Screenshot_2024-01-07_16-38-26.png
Screenshot_2024-01-07_16-38-26.png
  458   Fri Jan 5 12:29:10 2024 JM, CBFurther bias tests

Biased detectors this morning to 150V, the leakage current for all detectors appears stable and is ~3.5 uA. We have also replaced the CAEN bias module so that all channels are now functional. Detectors are connected as follows;

CHN0 -> Top left

CHN1 -> Top Right

CHN2 -> Bottom Left

CHN3 -> Bottom Right

Reconnected the motors and moved the detectors while biased. No observable change in the leakage current during movement at 150V. An increase in the pressure of the electron target was observed during movement (up to 1e-9 mbar). This increase was only momentarily, and the pressure recovered after a couple of seconds. We expect the pressure increase from movement to decrease over time.

 

  457   Thu Jan 4 13:15:10 2024 CB, JMTent removed - bias test

Removed tent, octagon, Neg cables.

Disconnected and removed Baking Pi, TC08, thermocouples.

Tested new rails, plates, fees - it fits. Need to check on the flange next.

Will soon turn off CARME IE514 gauge to bias test the DSSDs.

All four detectors were bias tested using the new flexi-rigid adaptor cards. Note the LK-1 jumper was removed from the adaptor cards so that 'open line' was observed between the bias lemo and the ground. Attachment 1 shows VI curves for all detectors. The maximum leakage current for all detectors appears lower than previous bias tests eg elog 400 -> temperature was ~17C today compared to ~20C for elog 400. We also performed a bias test using the old adaptor cards to see if there was any difference between the old and new cards. No significant difference was observed. For the top left detector we initally observed a significantly large leakage current (4.2 uA at 20V, 8.7 uA at 30V) which had ~1 uA fluctuations. We left for 5 minutes to see if it stabilised and the leakage current suddenly dropped to what is shown in the plot. This leakage current then appeared stable for the rest of the test. Perhaps this effect was observed by Tom but for the top right detector in elog 421?

We plan to connect the bias as we did in the 14N+d measurement so that we can monitor and plot the leakage current over the weekend.

Attachment 1: Bias.png
Bias.png
  456   Wed Oct 18 16:47:36 2023 JG, CBPower outage

Local power outage at the CRYRING today, while JG was trying to activate the ion pumps. Unlcear what caused it.

Valve on scroll worked as intended, but CARME reacehd 1E-7 mbar from turbo (?) venting.

Power restored. Pressure back in the 1E-11 mbar range, decreasing.

Temperature still well above room.

  455   Tue Oct 17 08:43:47 2023 CBNEG activation -2

09:40 - Found temperature and pressure stablised, but too high to comfortably activate NEGs closest to detectors. Detector temperature around 90C. This was expected. Ramped down UHV1400 #1,2,3,4 to P=50W (200C?). Ramped down heaters to 100, 80, 100C. Ramped down heating jacket to 60C.

10:00 - Activated set #5 (x2) at 80W tramp=30, thold=infinity; set#6 (x3) at 120W rest as #5. This is different from previous procedure because we have 4x DSSD but sensor only one one of them. It seems less dangerous to heat up everything at once so that the one TC we have is hopefully representative of the situation.

10:15 - CARME cooling. No visible effects on temperature. Tried P#5 100, P#6 150 W. (500C roughly?) Will probably have to decrease.

10:25 - Detectors getting hotter quickly, but still below 100C. No changes for now.

10:35 - As expected, detectors are getting too hot, too quickly. Decreased power to 60, 90 W

10:43 - Ramped down #5,6 to 40, 60 W. Should be around 200C at equilibrium.

10:45 - Also decreased #1,2,3,4 to 40W, and D2000 to 150C.

10:52 - Further decreased #5,6 to 10, 15W.

11:04 - Temperature back under control

11:10 - Detectors actually cooling now. Increased #5,6 to 20,30 W.

11:18 - Detectors still cooling. Increased #5.6 to 30,45W. If stable plan to leave like this until later this afternoon.

13:23 - Temperatures stable. Increased #5,6 to 40, 60 W. Decreased heaters to 80,80,80 C. Decreased jacket to 40C. Opened the tent partly to connect the vacuum division non-bakeable ion pump cable.

13:35 - Ion pump cable connected and tent closed as best I can. Interestingly, the pressure is going up probably as a result of the detector/kapton cables temperature going up due to the slight increase in #5,6 power.

14:15? - Graziano S. flashed the vacuum divison ion pump.

14:25 - Pressure keeps increasing slightly. I am taking this as a sign the kapton is outgassing, which is being made to out gass which is good. Decided to leave temperatures as they are until later today and start cooling down in the late afternoon / evening. Decreased slightly #5,6 to 34, 51W. Increased D2000 to 200C. Will leave like this if stable.

15:00 - Looks stable. See attachment 3. Leaving for a few hours, before proceeding to gradual cool down overnight. Will ask Jan for help in turning everything really off.

18:00 - All NEGs off. All apadev 10&11 channels to 40C. Ion pumps remain off. Asked JG to come tomorrow late morning and turn the apadev channels off and the ions pumps on.

Attachment 1: Screenshot_2023-10-17_09-45-23.png
Screenshot_2023-10-17_09-45-23.png
Attachment 2: Screenshot_2023-10-17_11-04-09.png
Screenshot_2023-10-17_11-04-09.png
Attachment 3: Screenshot_2023-10-17_14-56-55.png
Screenshot_2023-10-17_14-56-55.png
Attachment 4: Screenshot_2023-10-17_17-58-24.png
Screenshot_2023-10-17_17-58-24.png
Attachment 5: Screenshot_2023-10-17_17-58-51.png
Screenshot_2023-10-17_17-58-51.png
  454   Mon Oct 16 08:28:02 2023 CBNEG activation -1

9:00 - Michael's online vacuum pressure plot is no longer working - an old image is being displayed. CARME temperatures around 80-85C. Lowered heater temperatures to 150, 100, 150 C for safety. Lowered jacket temperature to 80C

9:15 - Flashed ion pumps several times. Small pressure increase, up to ~7E-9 but not macroscopic

9:30 - Started D-2000 #1,2. T=400, tramp 30, thold 90 (this was a mistake 1,3 are behind the detectors, 2,4 are to the side. They should ideally be activated together. Got confused. Not a big deal.)

9:50 - Online vacuum plot is working again. Pressure changing, internals around 80C

10:00 - Internal temperature still stable at 80C. Started D-2000 #3,4. T=400, tramp 30, thold 90

11:30 - For unclerar reasons the pressure has been increasing over the last half hour, after the four NEGs were activated. Nikos and the alignment group were adjusting the feet of the interaction chamber to improve the alignment. Hopefully this is not a leak. Section 10 was apparently vented overnight, and may be in air at the moment.

11:45 - D2000 activation complete. Set temperature to 200C for all four of them. Altough the temperature is still >375C, the pressure drops very quickly. This is good, but I have not observed this behaviour before.

13:10 - Pressure and temperature mostly stable. Turning on UHV1400 groups #1, 2 with at P=250W, tramp 30s, thold 90s.

13:35 - No effect on DSSD temperature. Increasing P->300W as per previous activation procedure.

13:43 - Still no sizeable effect. Increased P->350 W

15:00 - Completed activation of UHV1400 sets #1,2. Set power to 100W. Activated sets #3,4 at P=250W, same settings as other two.

15:30 - Increased UHV1400 #3,4 to P=300 W. Temperatures OK.

15:37 - Increased UHV1400 #3,4 to P=350 W. Temperatures increasing but still acceptable.

16:06 - Temperatures are starting to get a bit too high. Decreased P->300W. Also decreased heaters to 125, 80, 125C and baking jacket to 70 C.

16:26 - Decreased UHV1400 #3,4 to P=250W.

17:05 - Completed activation of UHV1400 sets #3,4. Set power to 100W. Planning to leave stabilising overnight.

Attachment 1: Screenshot_2023-10-16_17-17-30.png
Screenshot_2023-10-16_17-17-30.png
  453   Sun Oct 15 09:35:03 2023 CBStarting ramp down

10:40 - Status as attached. Heaters setpoint/read at 250 C, 150 C, 225 C. Power stable at 53%, 41%, 63%

10:45 - Informed HKR ramp down / pump activation is starting for CARME only. Setpoints to 200, 125, 200 C. Monitoring slow ramp down.

14:10 - Heater temperatures reached the set points. Power stabilising. CARME internals temperature dropped to around 100 C and still falling. Waiting until later today to further reduce heater setpoint. Will probably aim for 80 C before beginning activation.

15:45 - Heater setpoints to 175, 100, 175 C

15:50 - Interaction chamber jacket setpoints to 100C

18:00 -  CARME temperature decreasing, not stable yet. Will check later tonight.

20:43 - CARME internals around 85C, not quite stable still. Waiting overnight.

Attachment 1: Screenshot_2023-10-15_10-35-47.png
Screenshot_2023-10-15_10-35-47.png
Attachment 2: Screenshot_2023-10-15_17-56-00.png
Screenshot_2023-10-15_17-56-00.png
  452   Tue Oct 3 08:48:36 2023 TDTuesday 3 October
09.40 CARME IE514 1.79e-8 mbar, ambient temperature 20.7 deg C

      CARME bakeout monitor, ELOTECH R4000 temperature controller front panel and Grafana bakeout dashboard- attachments 1-3

      No change to ELOTECH R4000 set point
Attachment 1: 1000007432.jpg
1000007432.jpg
Attachment 2: 1000007431.jpg
1000007431.jpg
Attachment 3: 1000007434.jpg
1000007434.jpg
  451   Mon Oct 2 08:41:43 2023 TDMonday 2 October
09.35 CARME IE514 1.95e-8 mbar, ambient temperature 20.8 deg C

      CARME bakeout monitor, ELOTECH R4000 temperature controller front panel and Grafana bakeout dashboard- attachments 1-3

      No change to ELOTECH R4000 set point

17.45 CARME IE514 1.91e-8 mbar, ambient temperature 20.8 deg C

      CARME bakeout monitor, ELOTECH R4000 temperature controller front panel and Grafana bakeout dashboard- attachments 4-6

      No change to ELOTECH R4000 set point
Attachment 1: 1000007409.jpg
1000007409.jpg
Attachment 2: 1000007408.jpg
1000007408.jpg
Attachment 3: 1000007407.jpg
1000007407.jpg
Attachment 4: 1000007416.jpg
1000007416.jpg
Attachment 5: 1000007415.jpg
1000007415.jpg
Attachment 6: 1000007413.jpg
1000007413.jpg
  450   Sun Oct 1 10:24:50 2023 TDSunday 1 October
11.20 CARME IE514 3.53e-9 mbar, ambient temperature 20.6 deg C

      CARME bakeout monitor, ELOTECH R4000 temperature controller front panel and Grafana bakeout dashboard- attachments 1-3

      Change ELOTECH R4000 set points for CARME heaters #1 and #3 to 200 deg C, heater #2 to 150 deg C

14.10 CARME IE514 6.12e-9 mbar, ambient temperature 20.7 deg C

      CARME bakeout monitor, ELOTECH R4000 temperature controller front panel and Grafana bakeout dashboard- attachments 4-6

      Change ELOTECH R4000 set points for CARME heaters #1, #2 and #3 to 225, 175 and 225 deg C respectively.


18.10 CARME IE514 1.92e-8 mbar, ambient temperature 20.8 deg C

      CARME bakeout monitor, ELOTECH R4000 temperature controller front panel and Grafana bakeout dashboard- attachments 7-9

      Change ELOTECH R4000 set points for CARME heaters #1, #2 and #3 to 250, 150 and 225 deg C respectively.
      All other ( active ) ELOTECH R4000 set points ( interaction chamber etc ) to 130 deg C.

20.25 CARME IE514 2.14e-8 mbar, ambient temperature 20.8 deg C

      CARME bakeout monitor, ELOTECH R4000 temperature controller front panel and Grafana bakeout dashboard- attachments 10-12

      No change to ELOTECH R4000 set points
Attachment 1: 1000007352.jpg
1000007352.jpg
Attachment 2: 1000007354.jpg
1000007354.jpg
Attachment 3: 1000007353.jpg
1000007353.jpg
Attachment 4: 1000007360.jpg
1000007360.jpg
Attachment 5: 1000007358.jpg
1000007358.jpg
Attachment 6: 1000007361.jpg
1000007361.jpg
Attachment 7: 1000007368.jpg
1000007368.jpg
Attachment 8: 1000007369.jpg
1000007369.jpg
Attachment 9: 1000007370.jpg
1000007370.jpg
Attachment 10: 1000007377.jpg
1000007377.jpg
Attachment 11: 1000007375.jpg
1000007375.jpg
Attachment 12: 1000007374.jpg
1000007374.jpg
  449   Sat Sep 30 10:58:14 2023 TDSaturday 30 September
10.50 CARME IE514 7.14e-10 mbar, ambient temperature 20.8 deg C

      CARME bakeout monitor, ELOTECH R4000 temperature controller front panel and Grafana bakeout dashboard- attachments 1-3

      Change ELOTECH R4000 set points for CARME heaters #1 and #3 to 125 deg C

14.20 CARME IE514 8.25e-10 mbar, ambient temperature 20.8 deg C

      CARME bakeout monitor, ELOTECH R4000 temperature controller front panel and Grafana bakeout dashboard- attachments 4-6

      Change ELOTECH R4000 set points for CARME heaters #1 and #3 to 150 deg C

18.30 CARME IE514 1.29e-9 mbar, ambient temperature 20.7 deg C

      CARME bakeout monitor, ELOTECH R4000 temperature controller front panel and Grafana bakeout dashboard- attachments 7-9

      Change ELOTECH R4000 set points for CARME heaters #1 and #3 to 175 deg C and heater #2 to 125 deg C
Attachment 1: 1000007236.jpg
1000007236.jpg
Attachment 2: 1000007239.jpg
1000007239.jpg
Attachment 3: 1000007237.jpg
1000007237.jpg
Attachment 4: 1000007251.jpg
1000007251.jpg
Attachment 5: 1000007250.jpg
1000007250.jpg
Attachment 6: 1000007252.jpg
1000007252.jpg
Attachment 7: 1000007256.jpg
1000007256.jpg
Attachment 8: 1000007257.jpg
1000007257.jpg
Attachment 9: 1000007255.jpg
1000007255.jpg
  448   Fri Sep 29 08:54:29 2023 TDFriday 29 September
09.50 CARME IE514 9.0e-11 mbar, ambient temperature 21.3 deg C

      CARME bakeout monitor and ELOTECH temperature controller - attachments 1 & 2


10.15 Commence bakeout ramp
      Set all ELOTECH PID temperature controller zones ( channels ) to 50 deg C
      Expect all non-CARME zones to reach temperature set point quickly. There will be a delay for the TCs corresponding to the CARME heaters but the internal CAME TCs 
will respond more slowly- will observe response of CARME TCs over time.

10.35 ELOTECH temperature controller - attachment 3

11.10 CARME TCs change from c. 25-28 deg C to 28-35 deg C

11.15 Set all ELOTECH PID temperature controller zones ( channels ) to 100 deg C

12.15 CARME IE514 1.39e-10 mbar, ambient temperature 20.8 deg C

      CARME bakeout monitor and ELOTECH temperature controller - attachments 4 & 5


14.20 CARME IE514 3.62e-10 mbar, ambient temperature 20.8 deg C

      CARME bakeout monitor, ELOTECH temperature controller front panel and Grafana bakeout dashboard- attachments 6-8

15.20 Whilst all of the ELOTECH temperature controller TCs are at their current set point ( 100 deg C ), the CARME TCs are not yet at equilibrium. Plan to leave system 
in current state until tomorrow morning. Assuming CARME is at equilibrium will then increment ELOTECH temperature controller TC set points.

17.50 CARME IE514 6.59e-10 mbar, ambient temperature 20.6 deg C

      CARME bakeout monitor, ELOTECH temperature controller front panel and Grafana bakeout dashboard- attachments 9-11
Attachment 1: 1000007175.jpg
1000007175.jpg
Attachment 2: 1000007174.jpg
1000007174.jpg
Attachment 3: 1000007176.jpg
1000007176.jpg
Attachment 4: 1000007178.jpg
1000007178.jpg
Attachment 5: 1000007177.jpg
1000007177.jpg
Attachment 6: 1000007189.jpg
1000007189.jpg
Attachment 7: 1000007190.jpg
1000007190.jpg
Attachment 8: 1000007188.jpg
1000007188.jpg
Attachment 9: 1000007227.jpg
1000007227.jpg
Attachment 10: 1000007228.jpg
1000007228.jpg
Attachment 11: 1000007229.jpg
1000007229.jpg
  447   Thu Sep 28 09:19:50 2023 TDThursday 28 September
10.20 CARME IE514 8.6e-11 mbar

      CARME bakeout monitor and ELOTECH bakeout controller status - see attachments 1 & 2

      The CARME bakeout monitor shows recovery from test yesterday evening. The ELOTECH controller was switched to manual mode for the 3x CARME zones and 5% power selected 
for 45 minutes. CARME TCs increased from mid-20s to high 20s/low 30s over a period of 45 minutes. At the conclusion of the test power was set to 0% for all 3x CARME zones 
and the mode changed from manual to controlled.
Attachment 1: 1000007142.jpg
1000007142.jpg
Attachment 2: 1000007143.jpg
1000007143.jpg
  446   Wed Sep 27 15:56:41 2023 TDWednesday 27 September
16.55 CARME IE514 8.2e-11 mbar

      Vacuum pressure increased slowly overnight from c. 6e-11 mbar yesterday to c. 8e-11 mbar today following low power startup tests of the bakeout controller heating 
the CARME interaction and detector chambers yesterday and today.


Attachment 1 - bakeout controller config courtesy Anton Kalinin

Attachments 2 & 3 - bakeout controller summary displays and CARME bakeout monitor display c. 17.00 today

The 3x CARME heaters are connected to controller #4 zones # 1-3, i.e. lower right display, zones 1-3, mains fuses F9-F11

17.30 3x ( of 4x ) SAES D-2000 ion pump controllers ( SIP modules ) switched off. Fourth controller was already switched off and SHV cable disconnected from rear panel of 
SIP module.
Attachment 1: 1000007129.jpg
1000007129.jpg
Attachment 2: 1000007127.jpg
1000007127.jpg
Attachment 3: 1000007128.jpg
1000007128.jpg
  445   Tue Sep 26 10:13:14 2023 TDELOTECH PID Temperature Controller R4000 manual ( German and English )
Attachment 1: BA-Isoheat_R4000-XX-X-000-X-X_DE_V1.03.pdf
BA-Isoheat_R4000-XX-X-000-X-X_DE_V1.03.pdf BA-Isoheat_R4000-XX-X-000-X-X_DE_V1.03.pdf BA-Isoheat_R4000-XX-X-000-X-X_DE_V1.03.pdf BA-Isoheat_R4000-XX-X-000-X-X_DE_V1.03.pdf BA-Isoheat_R4000-XX-X-000-X-X_DE_V1.03.pdf BA-Isoheat_R4000-XX-X-000-X-X_DE_V1.03.pdf BA-Isoheat_R4000-XX-X-000-X-X_DE_V1.03.pdf BA-Isoheat_R4000-XX-X-000-X-X_DE_V1.03.pdf
Attachment 2: R4000-XX-X-000-X-X_EN_V1.30.pdf
R4000-XX-X-000-X-X_EN_V1.30.pdf R4000-XX-X-000-X-X_EN_V1.30.pdf R4000-XX-X-000-X-X_EN_V1.30.pdf R4000-XX-X-000-X-X_EN_V1.30.pdf R4000-XX-X-000-X-X_EN_V1.30.pdf R4000-XX-X-000-X-X_EN_V1.30.pdf R4000-XX-X-000-X-X_EN_V1.30.pdf R4000-XX-X-000-X-X_EN_V1.30.pdf
  444   Tue Sep 26 08:08:42 2023 TDTuesday 26 September
09.07 IE514 pressure 6.4e-11 mbar, ambient temperature +20.3 deg C

      Using appc218 ( account carme ) workstation by the CARME 19" rack 
     
      ssh pi@carmebakingpi

      Re-connect to bakeout monitor app 

      screen -dR

      See attcahment 1


12.02 Vacuum group ion pump disconnected and RGA removed from CARME chamber -  outside of CARME bakeout tent 'closed' - see attachments 2-4

      Vacuum group IE514 to be disconnected


14.45 Vacuum group disconnect their IE514

      Inside of CARME bakeout tent closed

      Photos of the CARME bakeout tent inside and outside CRYRING - see attachments 5-7  
Attachment 1: 1000007087.jpg
1000007087.jpg
Attachment 2: 1000007090.jpg
1000007090.jpg
Attachment 3: 1000007091.jpg
1000007091.jpg
Attachment 4: 1000007089.jpg
1000007089.jpg
Attachment 5: 1000007096.jpg
1000007096.jpg
Attachment 6: 1000007095.jpg
1000007095.jpg
Attachment 7: 1000007094.jpg
1000007094.jpg
  443   Mon Sep 25 16:33:32 2023 TDRemoved tape from heaters TC
17.33 IE514 pressure 6.4e-11 mbar, ambient temperature +20.8 deg C

      The thermocouples on the 3x groups of heater elements have been detached from direct contact with the heating element and now float abpove the heating element by 2-
3cm - see attachments 1 & 2

      The vacuum group IE514 and Ametek RGA are still attached to CARME - see attachment 3. Have requested vacuum group remove them.

      The 'air side' of the Be window on the interaction chamber is under vacuum provided by small Agilent scroll pump.

      
Attachment 1: 1000007072.jpg
1000007072.jpg
Attachment 2: 1000007067.jpg
1000007067.jpg
Attachment 3: 1000007070.jpg
1000007070.jpg
ELOG V3.1.4-unknown